OpenStudy (micheal1234):

Lesson Vocabulary Addition, Subtraction, Multiplication & Division Properties of Equality Equivalent Equations Isolate Inverse Operations 10. Two college friends rent an apartment. They have to pay the landlord two months’ rent and a $500 security deposit when they sign the lease. The total amount they pay the landlord is $2800. What is the rent for one month? 11. Solve for the variable: 20 – 3h = 2

OpenStudy (micheal1234):

please help

OpenStudy (anonymous):

We can model 10 as an equation Let t be one month's rent, therefore 2t is twice the month's rent \[2t+500=2800\]

OpenStudy (anonymous):

10. 2800-500=2 months rent (2300) 2300/2= 1 months rent ($1150) 11. To solve this, we have to isolate the variable. 20-3h=2 First, we can subtract 20 on both sides. This will give us -3h= -18 Then we can divide by -3 on both sides. This will give us h=6

OpenStudy (anonymous):

Would be better to try to let him do it on his own while we give hints, but I agree.

OpenStudy (micheal1234):

that what is though this was for